OpenStudy (anonymous):

the most efficient first step in the process to factor out the trinomial 2x3-18x2+40x is: A) Factor out -1 B) Factor out 2x C) Factor out (x-5) D) Factor out 2

OpenStudy (anonymous):

@hartnn @ganeshie8 can either of you help me?

hartnn (hartnn):

\(\Large 2x^3 = 2\times x\times x\times x \\ \Large -18x^2 = -1\times 2\times 3\times 3\times x\times x \\ \Large 40x = 2\times 2\times 2\times 5\times x\) what do you see common in these 3 terms?? that you can factor out :)

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Or is it just 2? @hartnn

OpenStudy (anonymous):

@perl is it 2x or just 2? can u help me?

OpenStudy (anonymous):

@MSxSammiBooxX

OpenStudy (anonymous):

@Squirrels can you help me?

hartnn (hartnn):

its 2x :)